OpenStudy (anonymous):

Jasson's Dessert for 4 serving 2 cups of skim milk 1 1/2 teaspoons of brown sugar How many teaspoons of brown sugar are needed for 6 cups of skim milk?

OpenStudy (anonymous):

How do i caculate it?? Can you show me step please?

OpenStudy (anonymous):

I thought people never ask :P

OpenStudy (anonymous):

Well i'm asking -_-

OpenStudy (anonymous):

x sugar/ 6 cups = 1.5/2 => x = 6 * 1.5 /2 = 4.5 cups of sugar.

OpenStudy (anonymous):

I means 4.5 teaspoons of brown sugar